# Training LLMs over Neurally Compressed Text

### v1

@@ 3.1 Training Data
All training data used is English web text from C4 (en 3.1.0) [52]. After tokenization, each document in C4 has an <EOS> token appended to it. We concatenate $128$ documents together to generate a long sequence of text. Using UTF-8 byte-level tokenization, the average document length is $2{,}170$ bytes, thus these long sequences have an average length of $277{,}760$ bytes. Despite the document breaks, we consider these long sequences “continguous” for the training of language models. These sequences are then split into individual examples, which are shuffled using the deterministic dataset functionality from SeqIO [54].
@@ 3.2 Training M1
The model used for compression is a decoder-only Transformer model [67]. It uses the $3$m size seen in Table 4 and a context length of $1{,}024$. We use a batch size of $128$, an rsqrt decay learning rate schedule ($1/\sqrt{steps}$) starting at $1.0$ with $10{,}000$ warmup steps, and a z-loss of $0.0001$. The model is trained for $2{,}500{,}000$ steps using the Adafactor [59] optimizer. The feed-forward layers use ReLU activations [47, 22], and we use distinct learnable relative attention embeddings [58] at each layer. We use a deterministic SeqIO dataset and train using Jax [7], Flax [28], and T5X [54]. The final validation performance of the M1 model is $1.457$ bits/byte, a standard measure of perplexity, see Section 3.8. M1 and M2 are both trained on the C4 training data, but the final validation data used to evaluate M2 is unseen during M1 training, therefore there is no information leakage. This is similar to how LLM tokenizers are often trained on same dataset that the LLM is subsequently trained on.
@@ 3.3 Compression Methods
When compressing C4 training data, we use an example length of $10{,}240$ bytes and apply one of the following compression techniques (see Appendix H for more methods we considered). This results in compressed examples that are, on average, much longer than our target sequence length of $512$ M2 tokens. Thus, each example fills or nearly fills the model’s context window with a compressed sequence made from contiguous raw bytes. We compress $51{,}200{,}000$ examples using each method, allowing us to train each M2 model for $200{,}000$ steps without repeating data.
@@
The compressor model is run over contiguous text sequences of $10{,}240$ bytes. The generated logits are used as the model distribution for Arithmetic Coding. We use the Range Encoding (a finite-precision implementation of Arithmetic Coding) implementation from TensorFlow Compression [4] with a precision of $14$. The range encoding implementation uses integers with precision+ $+\,2$ bits. This is enough to encode $16$-bit float logits, so should not cause numerical issues as our models are trained using bfloat16. While the compressor model is only trained on sequences of length $1{,}024$, it uses relative position embeddings in its attention layers. Thus, it can be applied to longer sequences. Some works observe decreased performance as inputs are scaled to lengths beyond those seen in training [66, 51], but we find that compression performance is similar in the two settings. Compressing sequences of length $1{,}024$ yields a compression ratio of $5.46$ while compressing sequences of length $10{,}240$ yields a ratio of $5.49$. This suggests the performance drop from long sequences has minimal effect on compression, or that the increased contextual information makes up this difference.
@@
In addition to resetting the AC encoder, we also reset the M1 model’s context. This means that each $W$ bits of output can be decoded independently, at the cost of a weaker M1 model due to the lack of context. As each window is fully self-contained, the model no longer has to learn to track Arithmetic Coding state variables over long distances.
In cases where “spare bits” are available at the end of a window (but not enough to add an additional symbol of text), we pad with zeros. This complicates the decoding algorithm, but the compression scheme remains lossless. See Appendix I for further discussion and an alternative padding approach that gives similar results.
@@
GZip: As a baseline, we also explore training over text compressed using GZip [17] as implemented in the Python [65] zlib library using the default compression level. GZip uses the DEFLATE algorithm—a combination of Huffman Trees [32] and LZ77 [77]. First LZ77 is used to replace repeated substrings in the text with pointers back to the original substring. Then a Huffman Tree is built for the current—LZ77 compressed—example and used to compress it. Note that this setting is dynamic, as the Huffman tree, and hence the binary codes for each character, are unique to the example. These experiments explore a setting where both the modeling and coding components of compression are different from Arithmetic Coding.
3.4 Tokenization of Compressed Text
@@
We convert the bitstream into a token sequence by grouping every $N$ bits into a token—resulting in a vocabulary size of $2^{N}$. We explore settings of $N\in$888816161616}, resulting in vocabulary sizes of $v\mathord{=}256$ and $v\mathord{=}65{,}536$. As the tokens are created from the compressed bitstream, we expect the distribution of tokens to be more uniform than the usual Zipfian [76] distribution of word or subword tokens, allowing us to use larger vocabularies without encountering issues of rare or unattested tokens.
@@ 3.5 Training M2 on Compressed Data
Each M2 model is trained for $200{,}000$ steps with a batch size of $256$ and a sequence length of $512$. Thus each model trains on $26.2$ billion tokens. Of these, the vast majority (over $98.9$%) are non-padding tokens; see Appendix C for details and Table 13 for the exact size of each dataset. As methods with higher compression ratios cover more raw text per token, we also include the total number of bytes in each dataset. Shuffling of training sets is seeded, and dataset state is checkpointed during training, so each training run results in the model seeing each example exactly once.
@@ 3.6 Baselines
Bytes
These baselines train directly over UTF-8 bytes, using the byte tokenizer from ByT5 [74]. The models see $26.2$ billion bytes total (see Table 13).
SentencePiece
These baselines train on text tokenized with the SentencePiece vocabulary of $32{,}000$ tokens from T5 [52]. The models see $112$ billion bytes total (see Table 13).
@@ 3.7 Numerical Stability
Methods like the rank-sorted algorithm used in LLMZip [63] may help alleviate these issues as only the order of tokens needs to match between settings. The development of alternate methods of LLM-based compression should keep numerical stability issues in mind and ideally alleviate these issues in the design of the algorithm. Increasing the level of quantization could also help reduce numerical noise issues, as differences would mostly be lost in quantization, but this would have a negative impact on the compression ratio.
@@ 3.8 Evaluation
We also compare models on how much computation (FLOPs) is required to perform inference over a given length of raw text (bytes). More specifically, we calculate M2’s expected FLOPs/byte by scaling FLOPs/token—approximated by $2\,\times\,\texttt{params}$ (excluding embedding parameters) following [35]—by the token-level compression rate (as tokens/byte). For methods using an M1 model during compression, the FLOPs/byte cost of M1 is added. [^10] While there is a computational cost to running GZip over the input text, we ignore it as it is insubstantial compared to the cost of running M2 model inference. For more details on the evaluation metrics see Appendix G.
We evaluate models on a sample of the C4 validation set. During evaluation, the model is run over $20$ batches or ~$2.6$ million tokens. These tokens represent different amounts of text based on the compression method, making it impractical to run evaluation on the same sequence of bytes for all methods. To confirm that our validation samples are large enough to be representative, for each method, we train five $25$m parameter models with different seeds. We find the final performance to be extremely stable, with the largest standard deviation in bits/byte being $0.0061$. Thus, the variance introduced from sampling the validation set is negligible. See Appendix B for more information about variance.
@@ 4 Results
Figure 3: Models trained over compressed text are compared against baseline models in terms of bits/byte ($\downarrow$) and inference FLOPs/bytes ($\downarrow$). The ArithmeticCoding and StaticAC settings are essentially unlearnable, with models failing to outperform naïve baselines (dashed lines) that assign equal probability to all tokens. EqualInfoAC and GZip outperform naïve baselines and show improvement with scale. EqualInfoAC is the strongest of the compression-based methods, with EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ outperforming the Bytes baseline at all sizes. While SentencePiece performs the best, the gap between EqualInfoAC and SentencePiece narrows with scale. See Appendix A for the exact values used in this and other graphs.
Simple methods of training over neural-compressed text fail
As seen in Fig. 3, the most obvious approach—compression using Arithmetic Coding with M1 assigning next-token probabilities—fails to learn anything. Regardless of scale, the model only learns to output a uniform distribution over tokens, the performance of which is denoted by the dashed line. As the Arithmetic Coding procedure is near optimal [45], the compression ratio is essentially determined by the loss of M1. Thus, even though the M2 model learns nothing useful, when scaled by the compression rate, this setting ends up with the same performance as the M1 model. Similarly, models trained over data compressed with StaticAC—where M1 is replaced with a static unigram model—fail to learn. This result suggests that the difficultly in learning stems from the complexity or brittleness of the Arithmetic Coding process itself, rather than from M2’s inability to model M1. Note that the weak “modeling” component of this compression scheme results in a much lower compression rate and thus worse bits/byte performance, despite the model also learning a uniform distribution.
@@
Our SentencePiece baseline outperforms all other methods, including our Bytes baseline, across all model sizes. On the surface, this result seems to run counter to the recent findings of [16], where their byte-level models outperformed subword (BPE) models at medium and large scales. The discrepancy is due to prioritizing different metrics. They report the model’s bit compression rate on fixed-length ($2{,}048$ byte) sequences. While this is one type of “fair” comparison, it disadvantages subword models, as they are trained to model dependencies longer than $2{,}048$ bytes (but never evaluated on this ability), and are allotted fewer inference FLOPs to process the same text, as compared to the byte-level models. Additionally, bit compression ratio penalizes subword models for having larger vocabulary sizes. By contrast, our evaluation tests what perplexity models achieve on sequences of the same length they were trained on, and compares models at matching FLOPs/byte cost. This aligns with our end goal, which is to train an LLM that achieves the best perplexity at whatever sequence length it can handle, given a fixed budget for training and inference.
@@
However, EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ outperforms the SentencePiece baseline in terms of tokens/byte. Models using EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ take fewer autoregressive steps to generate the same text than models using SentencePiece encoding. This has the potential to reduce generation latency, at the cost of reduced compute efficiency. This is a tradeoff that is often worth making in production. For instance, speculative decoding [42] is a popular approach that performs redundant computation in order to potentially accelerate auto-regressive steps.
@@
It is apparent from Fig. 3 that if FLOPs/byte were held constant, SentencePiece would achieve slightly better bits/byte than EqualInfoAC. However there is another axis along which EqualInfoAC may still be preferred. Setting aside inference FLOPs, all our SentencePiece models require $23$% longer sequences to encode the same text when compared to our best EqualInfoAC setting ($b\mathord{=}16$, $v\mathord{=}65$k). This means that regardless of FLOPs used, the SentencePiece models will take more decoder steps at inference time. It is up to the practitioner whether it is “worth it” to trade off some bits/byte performance in order to achieve shorter sequences. In many serving scenarios, decoder steps are a practical bottleneck for determining system latency, and there are cases where one may be willing to incur extra inference cost to reduce latency (e.g., speculative decoding [43]). To this end, it may be advantageous to scale up an EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ model to recover bits/byte performance while retaining the reduced latency. This can be seen visually in Fig. 4.
@@
Tokenizing compressed text using a larger $16$-bit vocabulary ($v\mathord{=}65$k) results in a 2×2$\times$2 × higher token compression rate, seen in the leftward shift of each curve in Fig. 6.111111The same trend holds for larger $64$ and $128$-bit windows, but the performance increase with scale is so slight that we omit them from the graph. See Table 10 for the exact values. For Arithmetic Coding methods, larger vocabulary also improves bits/byte, seen as a downward shift in the curves. However, for GZip, we see the opposite trend. Arithmetic Coding and GZip differ the most in their coding component, which suggests that the reason for this difference could lie there. Note that the header and footer present in GZip-compressed data do not explain this difference, see Appendix E. For EqualInfoAC$[b\mathord{=}16]$, moving from $v\mathord{=}256$ to $v\mathord{=}65$k results in each window corresponding to a single token, which increases the “stability” of the token $\rightarrow$ text mapping. This could be one reason for the performance gain; see Section 6.1 for more discussion of “stability”.
@@
Given the recent findings of [55], we anticipate that continuing to scale models beyond $2$ billion parameters is unlikely to deliver an “emergent” ability to learn over AC-compressed text, since the bits/byte metric we use is smooth.
@@
When scaling models, [30] recommend that training tokens should be scaled linearly with model size. However, in our experiments above, all models see the same number of tokens, regardless of model size. Consequently, our largest models may be somewhat “undertrained”. [^12] The undertraining of our $2$b models is also visible in their validation loss curves, which still have a significant decreasing slope at $200{,}000$ steps, showing the models have not yet converged. To test whether following the “scaling laws” recommendation influences our results, we reevaluate our models at earlier checkpoints selected to maintain a constant ratio of training data to model size. We find that all core trends are unchanged in this setting. See Appendix D for details.

### v2

@@ 3.1 Training Data
All training data used is English web text from C4 (en 3.1.0) (Raffel et al., 2020). After tokenization, each document in C4 has an <EOS> token appended to it. We concatenate $128$ documents together to generate a long sequence of text. Using UTF-8 byte-level tokenization, the average document length is $2{,}170$ bytes, thus these long sequences have an average length of $277{,}760$ bytes. Despite the document breaks, we consider these long sequences “continguous” for the training of language models. These sequences are then split into individual examples, which are shuffled using the deterministic dataset functionality from SeqIO (Roberts et al., 2023).
@@ 3.2 Training M1
The model used for compression is a decoder-only Transformer model (Vaswani et al., 2017). It uses the $3$m size seen in Table 4 and a context length of $1{,}024$. We use a batch size of $128$, an rsqrt decay learning rate schedule ($1/\sqrt{\text{steps}}$) starting at $1.0$ with $10{,}000$ warmup steps, and a z-loss of $0.0001$. The model is trained for $2{,}500{,}000$ steps using the Adafactor (Shazeer & Stern, 2018) optimizer. The feed-forward layers use ReLU activations (Nair & Hinton, 2010, Fukushima, 1975), and we use distinct learnable relative attention embeddings (Shaw et al., 2018) at each layer. We use a deterministic SeqIO dataset and train using Jax (Bradbury et al., 2018), Flax (Heek et al., 2020), and T5X (Roberts et al., 2023). The final validation performance of the M1 model is $1.457$ bits/byte, a standard measure of perplexity, see Section 3.8. M1 and M2 are both trained on the C4 training data, but the final validation data used to evaluate M2 is unseen during M1 training, therefore there is no information leakage. This is similar to how LLM tokenizers are often trained on same dataset that the LLM is subsequently trained on.
@@ 3.3 Compression Methods
When compressing C4 training data, we use an example length of $10{,}240$ bytes and apply one of the following compression techniques (see Appendix B for more methods we considered). This results in compressed examples that are, on average, much longer than our target sequence length of $512$ M2 tokens. Thus, each example fills or nearly fills the model’s context window with a compressed sequence made from contiguous raw bytes. We compress $51{,}200{,}000$ examples using each method, allowing us to train each M2 model for $200{,}000$ steps without repeating data.
@@
The compressor model is run over contiguous text sequences of $10{,}240$ bytes. The generated logits are used as the model distribution for Arithmetic Coding. We use the Range Encoding (a finite-precision implementation of Arithmetic Coding) implementation from TensorFlow Compression (Ballé et al., 2024) with a precision of $14$. The range encoding implementation uses integers with precision+ $+\,2$ bits. This is enough to encode $16$-bit float logits, so should not cause numerical issues as our models are trained using bfloat16. While the compressor model is only trained on sequences of length $1{,}024$, it uses relative position embeddings in its attention layers. Thus, it can be applied to longer sequences. Some works observe decreased performance as inputs are scaled to lengths beyond those seen in training (Varis & Bojar, 2021, Press et al., 2022), but we find that compression performance is similar in the two settings. Compressing sequences of length $1{,}024$ yields a compression ratio of $5.46$ while compressing sequences of length $10{,}240$ yields a ratio of $5.49$. This suggests the performance drop from long sequences has minimal effect on compression, or that the increased contextual information makes up this difference.
@@
In addition to resetting the AC encoder, we also reset the M1 model’s context. [^9] We investigate not resetting M1’s context in Section A.2 and find that the resets are important for good performance. This means that each $W$ bits of output can be decoded independently, at the cost of a weaker M1 model due to the lack of context. As each window is fully self-contained, the model no longer has to learn to track Arithmetic Coding state variables over long distances.
In cases where “spare bits” are available at the end of a window (but not enough to add an additional symbol of text), we pad with zeros. This complicates the decoding algorithm, but the compression scheme can remain lossless. See Section D.5 for further discussion, handling the end of the input sequence, and an alternative padding approach that gives similar results.
@@
GZip: As a baseline, we also explore training over text compressed using GZip (Deutsch, 1996) as implemented in the Python (Van Rossum & Drake, 2009) zlib library using the default compression level. GZip uses the DEFLATE algorithm—a combination of Huffman Trees (Huffman, 1952) and LZ77 (Ziv & Lempel, 1977). First LZ77 is used to replace repeated substrings in the text with pointers back to the original substring. Then a Huffman Tree is built for the current—LZ77 compressed—example and used to compress it. Note that this setting is dynamic, as the Huffman tree, and hence the binary codes for each character, are unique to the example. These experiments explore a setting where both the modeling and coding components of compression are different from Arithmetic Coding.
@@ 3.4 Tokenization of Compressed Text
We convert the bitstream into a token sequence by grouping every $N$ bits into a token—resulting in a vocabulary size of $2^{N}$. We explore settings of $N\in$888816161616}, resulting in vocabulary sizes of $v\mathord{=}256$ and $v\mathord{=}65{,}536$. As the tokens are created from the compressed bitstream, we expect the distribution of tokens to be more uniform than the usual Zipfian (Zipf, 1935) distribution of word or subword tokens, allowing us to use larger vocabularies without encountering issues of rare or unattested tokens.
@@ 3.5 Training M2 on Compressed Data
Each M2 model is trained for $200{,}000$ steps with a batch size of $256$ and a sequence length of $512$. Thus each model trains on $26.2$ billion tokens. Of these, the vast majority (over $98.9$%) are non-padding tokens; see Section D.2 for details and Table 12 for the exact size of each dataset. As methods with higher compression ratios cover more raw text per token, we also include the total number of bytes in each dataset. Shuffling of training sets is seeded, and dataset state is checkpointed during training, so each training run results in the model seeing each example exactly once.
@@ 3.6 Baselines
Bytes: These baselines train directly over UTF-8 bytes, using the byte tokenizer from ByT5 (Xue et al., 2022). The models see $26.2$ billion bytes total (see Table 12).
SentencePiece: These baselines train on text tokenized with the SentencePiece vocabulary of $32{,}000$ tokens from T5 (Raffel et al., 2020). The models see $112$ billion bytes total (see Table 12).
@@ 3.7 Numerical Stability
Methods like the rank-sorted algorithm used in LLMZip (Valmeekam et al., 2023) may help alleviate these issues as only the order of tokens needs to match between settings. The development of alternate methods of LLM-based compression should keep numerical stability issues in mind and ideally alleviate these issues in the design of the algorithm. Increasing the level of quantization could also help reduce numerical noise issues, as differences would mostly be lost in quantization, but this would have a negative impact on the compression ratio.
@@ 3.8 Evaluation
We also compare models on how much computation (FLOPs) is required to perform inference over a given length of raw text (bytes). More specifically, we calculate M2’s expected FLOPs/byte by scaling FLOPs/token—approximated by $2\,\times\,\texttt{params}$ (excluding embedding parameters) following (Kaplan et al., 2020)—by the token-level compression rate (as tokens/byte). For methods using an M1 model during compression, the FLOPs/byte cost of M1 is added. [^10] While there is a computational cost to running GZip over the input text, we ignore it as it is insubstantial compared to the cost of running M2 model inference. For more details on the evaluation metrics see Section D.4.
We evaluate models on a sample of the C4 validation set. During evaluation, the model is run over $20$ batches or ~$2.6$ million tokens. These tokens represent different amounts of text based on the compression method, making it impractical to run evaluation on the same sequence of bytes for all methods. To confirm that our validation samples are large enough to be representative, for each method, we train five $25$m parameter models with different seeds. We find the final performance to be extremely stable, with the largest standard deviation in bits/byte being $0.0061$. Thus, the variance introduced from sampling the validation set is negligible. See Section D.1 for more information about variance.
@@ 4 Results
Figure 3: Models trained over compressed text are compared against baseline models in terms of bits/byte ($\downarrow$) and inference FLOPs/bytes ($\downarrow$). The ArithmeticCoding and StaticAC settings are essentially unlearnable, with models failing to outperform naïve baselines (dashed lines) that assign equal probability to all tokens. EqualInfoAC and GZip outperform naïve baselines and show improvement with scale. EqualInfoAC is the strongest of the compression-based methods, with EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ outperforming the Bytes baseline at all sizes. While SentencePiece performs the best, the gap between EqualInfoAC and SentencePiece narrows with scale. See Section D.9 for the exact values used in this and other graphs.
Simple methods of training over neurally compressed text fail
As seen in Fig. 3, the most obvious approach—compression using Arithmetic Coding with M1 assigning next-token probabilities—fails to learn anything. Regardless of scale, the model only learns to output a uniform distribution over tokens, the performance of which is denoted by the dashed line. As the Arithmetic Coding procedure is near optimal (Mahoney, 2013), the compression ratio is essentially determined by the loss of M1. Thus, even though the M2 model learns nothing useful, when scaled by the compression rate, this setting ends up with the same performance as the M1 model. Similarly, models trained over data compressed with StaticAC—where M1 is replaced with a static unigram model—fail to learn. This result suggests that the difficultly in learning stems from the complexity or brittleness of the Arithmetic Coding process itself, rather than from M2’s inability to model M1. Note that the weak “modeling” component of this compression scheme results in a much lower compression rate and thus worse bits/byte performance, despite the model also learning a uniform distribution.
@@
Our SentencePiece baseline outperforms all other methods, including our Bytes baseline, across all model sizes. On the surface, this result seems to run counter to the recent findings of (Delétang et al., 2024), where their byte-level models outperformed subword (BPE) models at medium and large scales. The discrepancy is due to prioritizing different metrics. They report the model’s bit compression rate on fixed-length ($2{,}048$ byte) sequences. While this is one type of “fair” comparison, it disadvantages subword models, as they are trained to model dependencies longer than $2{,}048$ bytes (but never evaluated on this ability), and are allotted fewer inference FLOPs to process the same text, as compared to the byte-level models. Additionally, bit compression ratio penalizes subword models for having larger vocabulary sizes. By contrast, our evaluation tests what perplexity models achieve on sequences of the same length they were trained on, and compares models at matching FLOPs/byte cost. This aligns with our end goal, which is to train an LLM that achieves the best perplexity at whatever sequence length it can handle, given a fixed budget for training and inference.
@@
However, EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ outperforms the SentencePiece baseline in terms of tokens/byte. Models using EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ take fewer autoregressive steps to generate the same text than models using SentencePiece encoding. This has the potential to reduce generation latency, at the cost of reduced compute efficiency. This is a tradeoff that is often worth making in production. For instance, speculative decoding (Leviathan et al., 2023b) is a popular approach that performs redundant computation in order to potentially accelerate auto-regressive steps.
@@
It is apparent from Fig. 3 that if FLOPs/byte were held constant, SentencePiece would achieve slightly better bits/byte than EqualInfoAC. However there is another axis along which EqualInfoAC may still be preferred. Setting aside inference FLOPs, all our SentencePiece models require $23$% longer sequences to encode the same text when compared to our best EqualInfoAC setting ($b\mathord{=}16$, $v\mathord{=}65$k). This means that regardless of FLOPs used, the SentencePiece models will take more decoder steps at inference time. It is up to the practitioner whether it is “worth it” to trade off some bits/byte performance in order to achieve shorter sequences. In many serving scenarios, decoder steps are a practical bottleneck for determining system latency, and there are cases where one may be willing to incur extra inference cost to reduce latency (e.g., speculative decoding (Leviathan et al., 2023a)). To this end, it may be advantageous to scale up an EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$ model to recover bits/byte performance while retaining the reduced latency. This can be seen visually in Fig. 4.
@@
Tokenizing compressed text using a larger $16$-bit vocabulary ($v\mathord{=}65$k) results in a 2×2$\times$2 × higher token compression rate, seen in the leftward shift of each curve in Fig. 6.111111The same trend holds for larger $64$ and $128$-bit windows, but the performance increase with scale is so slight that we omit them from the graph. See Table 16 for the exact values. For Arithmetic Coding methods, larger vocabulary also improves bits/byte, seen as a downward shift in the curves. However, for GZip, we see the opposite trend. Arithmetic Coding and GZip differ the most in their coding component, which suggests that the reason for this difference could lie there. Note that the header and footer present in GZip-compressed data do not explain this difference, see Section A.4. For EqualInfoAC$[b\mathord{=}16]$, moving from $v\mathord{=}256$ to $v\mathord{=}65$k results in each window corresponding to a single token, which increases the “stability” of the token $\rightarrow$ text mapping. This could be one reason for the performance gain; see Section 6.1 for more discussion of “stability”.
@@
Given the recent findings of (Schaeffer et al., 2023), we anticipate that continuing to scale models beyond $2$ billion parameters is unlikely to deliver an “emergent” ability to learn over AC-compressed text, since the bits/byte metric we use is smooth.
@@
When scaling models, (Hoffmann et al., 2022) recommend that training tokens should be scaled linearly with model size. However, in our experiments above, all models see the same number of tokens, regardless of model size. Consequently, our largest models may be somewhat “undertrained”. [^12] The undertraining of our $2$b models is also visible in their validation loss curves, which still have a significant decreasing slope at $200{,}000$ steps, showing the models have not yet converged. To test whether following the “scaling laws” recommendation influences our results, we reevaluate our models at earlier checkpoints selected to maintain a constant ratio of training data to model size. We find that all core trends are unchanged in this setting. See Section D.3 for details.